OpenStudy (anonymous):

x^2 = 7x – 9 Solve using the quadratic formula. please show work

OpenStudy (anonymous):

x^2 - 7x + 9 = 0 x = [--7 +- sqrt(49 - 4 * 1 * 9) ] / 2 =[ 7 + sqrt 13] / 2 or [ 7 - sqrt13] / 2 = 3.5 + 0.5sqrt13 , 3.5 - 0.5sqrt13

OpenStudy (anonymous):

x = 5.303, 1.697 correct to 3 dec places

OpenStudy (anonymous):

ok happy with that

hero (hero):

Post your problem to the left

hero (hero):

x^2 – 6x = 16 x^2-6x - 16 = 0 x^2-8x+2x - 16 = 0 x(x-8)+2(x-8) = 0 (x-8)(x+2) = 0 x = -2,8 This one can be solved if you factor by grouping. No need to use quad formula.